Shift report 20020122 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 22 2002 - 23:58:52 EST

  • Next message: Apache: "Shift report 20020123 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Jan 22 23:58:52 2002, Shift: 16:00-24:00
    Supervisor: BHS, with: FV
    Log book#12 pages 106 - 106
    
    Shift Summary:
    ------------------------------------------------------------
    Beam studies until 1900, and then they took two more hours
    getting the beam up. Since then we have had a continous
    store.
    
    One data run:
    6740 - still going
    MRS 90, 6B
    FS 12, 1/2 A
    Flemming says to go on with this setting trough the night.
    
    I had problems with D2 - it would not turn on, and a mag-
    guy had to fiddle with it for 20mins before it would wake
    up. 
    TPC pedestsal run: 
    6738 - installed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 22 2002 - 23:59:25 EST